co. is now considering which of two product labeling machine to buy. Machine A
costs $5,000 to buy and $300 per year to operate. It wears out and must be replaced every two
years. Machine B costs $6,500 to buy and $200 per year to operate. It lasts for four years.
Ignoring taxes, which one should it choose, if it use a 10% discount rate?

i. Compute PV cost of each machine.
ii. Compute equivalent annual cost for each machine, and decide which to buy.
Answer #1
i] Machine A:
PV of cost = -5000-300*(1.1^2-1)/(0.1*1.1^2) = $     -5,520.66
Machine B:
PV of cost = -6500-200*(1.1^4-1)/(0.1*1.1^4) = $     -7,133.97
ii] Machine A:
Equivalent annual cost = -5520.66*0.1*1.1^2/(1.1^2-1) = $     -3,180.95
Machine B:
Equivalent annual cost = -7133.97*0.1*1.1^4/(1.1^4-1) = $     -2,250.56
DECISION:
Per equivalent annual cost, Machine B is to be bought
as it has lower equivalent annual cost.